# Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A

### v3

@@ Appendix F Dataset-level results
Falcon 7B 25.7 50.5 0/2 50.2 0/2
Falcon 40B 53.0 66.8 2/2 59.8 2/2
Llama 2 7B 45.2 62.8 2/2 59.8 2/2
Llama 2 70B 74.2 77.4 2/2 69.5 2/2
Llama 3.0 8B 74.7 81.8 2/2 78.7 2/2
Llama 3.0 70B 92.3 89.7 2/2 82.2 2/2
Llama 3.1 8B 77.4 82.2 2/2 77.6 2/2
Llama 3.1 70B 93.4 91.3 2/2 78.6 2/2
Mistral 7B 70.7 67.3 2/2 67.3 2/2
Mixtral 8x7B 84.2 63.5 2/2 64.3 2/2
SOLAR 10.7B 80.5 61.4 2/2 71.2 2/2
Yi 6B 68.8 75.6 2/2 64.1 2/2
Yi 34B 85.4 67.3 2/2 70.0 2/2
GPT-3.5 Turbo 83.3 84.4 2/2 2/2
GPT-4o 96.3 88.4 2/2 2/2
Table 8: AUROC results for ARC-Challenge. See Table 1 for more explanation.
@@
Falcon 7B 24.9 50.1 0/2 49.4 0/2
Falcon 40B 42.8 61.8 2/2 59.2 2/2
Llama 2 7B 41.5 57.3 2/2 53.8 1/2
Llama 2 70B 63.9 70.0 2/2 64.7 2/2
Llama 3.0 8B 65.6 73.5 2/2 70.8 2/2
Llama 3.0 70B 77.6 81.6 2/2 70.4 2/2
Llama 3.1 8B 66.1 72.9 2/2 67.5 2/2
Llama 3.1 70B 78.9 81.7 2/2 59.8 2/2
Mistral 7B 52.5 63.0 2/2 62.6 2/2
Mixtral 8x7B 65.8 57.9 2/2 59.0 2/2
SOLAR 10.7B 79.2 63.5 2/2 69.0 2/2
Yi 6B 42.0 67.0 2/2 60.6 2/2
Yi 34B 75.4 71.8 2/2 68.7 2/2
GPT-3.5 Turbo 72.7 77.0 2/2 2/2
GPT-4o 88.4 87.9 2/2 2/2
Table 9: AUROC results for HellaSwag. See Table 1 for more explanation.
@@
Falcon 7B 25.8 50.1 0/2 50.9 0/2
Falcon 40B 43.9 62.0 2/2 52.9 1/2
Llama 2 7B 40.2 61.7 2/2 59.9 2/2
Llama 2 70B 56.0 71.8 2/2 65.9 2/2
Llama 3.0 8B 56.3 76.2 2/2 74.5 2/2
Llama 3.0 70B 75.0 83.5 2/2 78.6 2/2
Llama 3.1 8B 58.1 78.1 2/2 73.9 2/2
Llama 3.1 70B 79.8 84.1 2/2 72.9 2/2
Mistral 7B 52.5 64.9 2/2 64.8 2/2
Mixtral 8x7B 65.2 60.6 2/2 64.0 2/2
SOLAR 10.7B 58.0 60.8 2/2 66.9 2/2
Yi 6B 51.9 69.8 2/2 62.3 2/2
Yi 34B 65.6 64.1 2/2 65.7 2/2
GPT-3.5 Turbo 65.4 79.8 2/2 2/2
GPT-4o 84.0 85.2 2/2 2/2
Table 10: AUROC results for MMLU. See Table 1 for more explanation.
@@
Falcon 7B 20.5 62.4 2/2 55.8 0/2
Falcon 40B 28.7 53.4 0/2 50.6 0/2
Llama 2 7B 24.1 52.2 0/2 53.9 0/2
Llama 2 70B 44.6 71.5 2/2 66.9 2/2
Llama 3.0 8B 40.4 68.0 2/2 64.0 2/2
Llama 3.0 70B 72.1 79.8 2/2 70.9 2/2
Llama 3.1 8B 55.3 72.8 2/2 60.8 2/2
Llama 3.1 70B 74.1 83.6 2/2 66.2 2/2
Mistral 7B 55.1 66.0 2/2 63.4 2/2
Mixtral 8x7B 67.3 65.6 2/2 64.5 2/2
SOLAR 10.7B 49.1 58.1 1/2 63.1 2/2
Yi 6B 43.7 64.5 2/2 60.6 2/2
Yi 34B 56.7 66.8 2/2 67.8 2/2
GPT-3.5 Turbo 55.4 74.8 2/2 2/2
GPT-4o 82.4 81.3 2/2 2/2
Table 11: AUROC results for TruthfulQA. See Table 1 for more explanation.

### v4

@@ A.3 Details on base models
Falcon 7B 30.2 52.1 1/10 51.0 0/10
Falcon 40B 35.2 58.8 7/10 53.0 2/10
Llama 2 7B 32.1 54.6 3/10 52.2 1/10
Llama 2 70B 51.7 71.2 10/10 63.7 8/10
Llama 3.0 8B 48.0 65.4 8/10 58.2 6/10
Llama 3.0 70B 62.9 73.8 10/10 68.5 10/10
Llama 3.1 8B 47.2 65.2 9/10 60.6 7/10
Llama 3.1 70B 63.4 72.3 9/10 68.2 10/10
Mistral 7B 48.7 65.4 7/10 61.1 8/10
Mixtral 8x7B 52.1 68.1 10/10 61.9 8/10
SOLAR 10.7B 68.0 77.3 10/10 69.5 10/10
Yi 6B 37.2 60.6 9/10 62.5 10/10
Yi 34B 44.1 71.1 10/10 69.6 10/10
Table 9: Q&A-with-abstention results for the base models. We see roughly the same patterns as in the Q&A-with-abstention results for the chat models (Figure˜9). See Table 4 for an explanation of the scoring scheme.
Balanced Conservative
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 39.6-39.6 $0.1$ 0.4-0.4 109.3-109.3 0.5-0.5 1.3-1.3
Falcon 40B 29.6-29.6 $0.1$ 1.1-1.1 94.4-94.4 0.4-0.4 $0.0$
Llama 2 7B 35.7-35.7 2.7-2.7 0.1-0.1 103.5-103.5 0.7-0.7 0.3-0.3
Llama 2 70B $3.4$ $15.6$ $12.1$ 44.9-44.9 $8.1$ $5.2$
Llama 3.0 8B 4.0-4.0 $8.5$ $0.3$ 56.0-56.0 $5.3$ $1.1$
Llama 3.0 70B $25.9$ $33.3$ $30.8$ 11.2-11.2 $23.0$ $15.6$
Llama 3.1 8B 5.7-5.7 $9.3$ $5.6$ 58.5-58.5 3.8-3.8 $1.3$
Llama 3.1 70B $26.6$ $26.6$ $30.9$ 10.1-10.1 $4.5$ $4.0$
Mistral 7B 2.5-2.5 $10.7$ $6.7$ 53.7-53.7 $3.3$ $0.2$
Mixtral 8x7B $4.2$ $15.3$ $9.2$ 43.6-43.6 $9.0$ $3.0$
SOLAR 10.7B $35.9$ $40.9$ $36.0$ $3.8$ $22.5$ $15.8$
Yi 6B 25.7-25.7 $3.2$ $1.3$ 88.5-88.5 0.9-0.9 0.4-0.4
Yi 34B 11.7-11.7 $10.2$ $11.4$ 67.5-67.5 $6.2$ $9.2$
Appendix B Results for 5-shot prompting
As discussed in Section˜3, our main experiments used zero-shot prompting in order to test our hypothesis in the simplest setting possible. In this section, we show how the results change if 5-shot prompting is used instead. In short, the same correlations and trends hold for MSP (although not Max Logit). The AUROC values are also slightly lower overall, which could suggest that few-shot prompting disrupts the innate uncertainty information in LLMs, but we do not have the evidence to conclude this definitively. Our 5-shot experiments used the same setup as our zero-shot experiments, simply with 5 examples prepended to each prompt.
B.1 Chat models
Figure 11: Calibration results for 5-shot prompting for chat models. Calibration is somewhat worse than for zero-shot prompting with chat models (Figure˜2), but this may be noise. Similar to zero-shot prompting, there is no statistical evidence for a correlation between calibration error and Q&A accuracy ($R^{2}=0.03,p=0.57$).
Figure 12: AUROC results for 5-shot prompting for chat models. AUROC values are worse compared to zero-shot, but this may be noise. MSP AUROC (left) retains a strong correlation with Q&A accuracy ($R^{2}=0.88,p<10^{-4}$), but the correlation with Max Logit (right) is much weaker here ($R^{2}=0.28,p=0.06$).
Table 10: AUROC results for 5-shot prompting for chat models. See Table 2 for an explanation of the $p$-values.
@@ B.1 Chat models
Falcon 7B 29.6 52.0 1/10 51.2 0/10
Falcon 40B 37.9 61.5 10/10 58.5 8/10
Llama 2 7B 29.8 50.8 0/10 50.7 0/10
Llama 2 70B 45.6 61.1 8/10 59.8 8/10
Llama 3.0 8B 40.7 58.5 8/10 59.4 7/10
Llama 3.0 70B 65.4 70.0 10/10 58.2 7/10
Llama 3.1 8B 51.4 65.3 10/10 61.3 10/10
Llama 3.1 70B 67.7 71.2 10/10 58.2 7/10
Mistral 7B 48.0 61.7 10/10 60.2 9/10
Mixtral 8x7B 57.6 63.1 10/10 54.5 4/10
SOLAR 10.7B 58.4 66.5 10/10 62.0 10/10
Yi 6B 39.2 61.6 9/10 57.8 6/10
Yi 34B 54.8 65.2 10/10 62.0 10/10
GPT-3.5 Turbo 58.8 69.4 10/10
GPT-4o 73.2 71.2 10/10
Table 11: Q&A-with-abstention results for 5-shot prompting for chat models. See Table 4 for an explanation of the scoring scheme.
Balanced Conservative
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 40.7-40.7 1.8-1.8 0.9-0.9 111.1-111.1 1.3-1.3 2.6-2.6
Falcon 40B 24.3-24.3 $1.4$ $0.3$ 86.5-86.5 1.2-1.2 0.2-0.2
Llama 2 7B 40.4-40.4 0.7-0.7 $0.0$ 110.6-110.6 1.7-1.7 1.1-1.1
Llama 2 70B 8.7-8.7 $4.0$ $2.3$ 63.1-63.1 $0.0$ $0.0$
Llama 3.0 8B 18.7-18.7 $0.6$ $0.9$ 78.0-78.0 0.9-0.9 0.1-0.1
Llama 3.0 70B $30.8$ $33.4$ $31.4$ 3.8-3.8 $15.6$ $6.1$
Llama 3.1 8B $2.8$ $11.2$ $8.6$ 45.8-45.8 $1.1$ $1.2$
Llama 3.1 70B $35.5$ $37.5$ $35.7$ $3.3$ $18.3$ $12.6$
Mistral 7B 4.0-4.0 $2.6$ $1.8$ 56.0-56.0 $0.0$ $0.1$
Mixtral 8x7B $15.1$ $15.2$ $15.9$ 27.3-27.3 $1.4$ 4.1-4.1
SOLAR 10.7B $16.9$ $21.3$ $21.2$ 24.7-24.7 $0.0$ $1.6$
Yi 6B 21.6-21.6 $2.4$ $0.5$ 82.4-82.4 $0.4$ 0.5-0.5
Yi 34B $9.6$ $9.7$ $12.7$ 35.6-35.6 $2.6$ $3.1$
GPT-3.5 Turbo $17.6$ $20.5$ 23.5-23.5 $10.6$
GPT-4o $46.6$ $47.8$ $19.9$ $28.5$
B.2 Base models
Figure 13: Calibration results for 5-shot prompting for base models. The results are similar to the zero-shot prompting results for base models, although the correlation between calibration error and Q&A accuracy is weaker here ($R^{2}=0.32,p=0.04$).
Table 12: AUROC results for 5-shot prompting for base models. See Table 2 for an explanation of the $p$-values.
@@ B.2 Base models
Falcon 7B 29.2 52.4 2/10 51.8 1/10
Falcon 40B 33.9 57.5 8/10 57.1 5/10
Llama 2 7B 35.3 57.3 6/10 53.2 4/10
Llama 2 70B 51.4 65.9 9/10 63.8 9/10
Llama 3.0 8B 45.0 62.7 6/10 59.3 6/10
Llama 3.0 70B 53.4 66.3 10/10 63.0 10/10
Llama 3.1 8B 46.6 65.4 9/10 60.9 8/10
Llama 3.1 70B 49.6 64.5 10/10 61.9 10/10
Mistral 7B 46.0 62.5 8/10 58.3 6/10
Mixtral 8x7B 51.1 64.5 10/10 56.2 6/10
SOLAR 10.7B 60.6 70.8 10/10 64.4 10/10
Yi 6B 29.4 59.2 8/10 53.9 4/10
Yi 34B 35.6 65.5 10/10 60.7 9/10
Figure 14: AUROC results for 5-shot prompting for base models. The results are similar to zero-shot results for base models. Here the coefficients of determination with Q&A accuracy are $R^{2}=0.74$ and $R^{2}=0.64$ for MSP (left) and Max Logit (right) AUROC respectively, both with $p<10^{-3}$.
Table 13: Q&A-with-abstention results for 5-shot prompting for base models. See Table 4 for an explanation of the scoring scheme.
Balanced Conservative
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 41.8-41.8 $0.0$ 3.2-3.2 112.6-112.6 1.3-1.3 0.2-0.2
Falcon 40B 32.3-32.3 1.8-1.8 4.5-4.5 98.5-98.5 7.5-7.5 2.4-2.4
Llama 2 7B 29.4-29.4 $0.2$ $0.0$ 94.0-94.0 0.7-0.7 0.2-0.2
Llama 2 70B $2.7$ $12.7$ $11.4$ 45.9-45.9 $3.9$ 1.3-1.3
Llama 3.0 8B 10.1-10.1 $7.2$ $4.4$ 65.1-65.1 2.7-2.7 $0.1$
Llama 3.0 70B $6.9$ $17.7$ $19.4$ 39.7-39.7 $7.5$ $5.9$
Llama 3.1 8B 6.8-6.8 $8.0$ $4.7$ 60.2-60.2 $1.7$ 1.9-1.9
Llama 3.1 70B 0.7-0.7 $14.4$ $10.5$ 51.1-51.1 $4.0$ $1.4$
Mistral 7B 8.1-8.1 $5.3$ $3.2$ 62.1-62.1 $0.8$ $0.1$
Mixtral 8x7B $2.3$ $11.7$ $7.6$ 46.6-46.6 $1.5$ $1.2$
SOLAR 10.7B $21.1$ $29.7$ $23.9$ 18.4-18.4 $18.0$ $9.5$
Yi 6B 41.2-41.2 $0.1$ 0.6-0.6 111.9-111.9 0.8-0.8 1.8-1.8
Yi 34B 28.8-28.8 $1.8$ $0.8$ 93.3-93.3 3.7-3.7 1.1-1.1
Appendix C Other measures of uncertainty
Up until this point, we have only considered the maximum softmax probability and the maximum logit. However, there are other ways to measure the uncertainty of a distribution. Here we consider margin (the difference between the maximum probability/logit and the second largest probability/logit) and entropy (only for probabilities). All results in this section use chat models and zero-shot prompting. We omit calibration analysis for the margin and entropy metrics, because calibration relates to the probability of a particular outcome, and margin and entropy do not correspond to probabilities of any outcome.
Table 14: AUROC results using the entropy of the softmax probabilities instead of the MSP. See Table 2 for explanation of the $p$-values.
Entropy
LLM Q&A accuracy AUROC $p<10^{-4}$
Falcon 7B 30.1 52.8 1/10
Falcon 40B 40.5 60.6 7/10
Llama 2 7B 39.7 58.9 6/10
Llama 2 70B 58.7 71.1 10/10
Llama 3.0 8B 59.1 71.0 10/10
Llama 3.0 70B 78.8 81.3 10/10
Llama 3.1 8B 62.0 72.3 10/10
Llama 3.1 70B 80.6 83.3 10/10
Mistral 7B 57.0 66.1 10/10
Mixtral 8x7B 69.9 70.0 10/10
SOLAR 10.7B 67.5 71.8 10/10
Yi 6B 49.6 68.5 10/10
Yi 34B 69.1 75.9 10/10
GPT-3.5 Turbo 67.5 76.2 10/10
GPT-4o 86.9 85.3 10/10
Figure 15: AUROC results using the entropy of the probability distribution as the uncertainty metric. The AUROC values are similar to those obtained from the MSP, and the entropy AUROC also exhibits a strong correlation with Q&A accuracy ($R^{2}=0.93,p<10^{-4}$).
Table 15: AUROC results using the margin of softmax probabilities and margin of logits instead of the MSP and maximum logit, respectively. See Table 2 for explanation of the $p$-values.
Probability margin Logit margin
LLM Q&A accuracy AUROC $p<10^{-4}$ AUROC $p<10^{-4}$
Falcon 7B 30.1 50.1 0/10 49.6 0/10
Falcon 40B 40.5 59.2 5/10 58.7 5/10
Llama 2 7B 39.7 58.3 6/10 58.1 6/10
Llama 2 70B 58.7 71.0 10/10 71.0 10/10
Llama 3.0 8B 59.1 71.1 10/10 70.9 10/10
Llama 3.0 70B 78.8 81.3 10/10 81.3 10/10
Llama 3.1 8B 62.0 72.9 10/10 72.7 10/10
Llama 3.1 70B 80.6 83.4 10/10 83.4 10/10
Mistral 7B 57.0 66.2 10/10 66.2 10/10
Mixtral 8x7B 69.9 70.0 10/10 70.1 10/10
SOLAR 10.7B 67.5 71.8 10/10 71.9 10/10
Yi 6B 49.6 66.0 10/10 65.4 10/10
Yi 34B 69.1 75.6 10/10 75.5 10/10
GPT-3.5 Turbo 67.5 76.0 10/10
GPT-4o 86.9 85.3 10/10
Figure 16: AUROC results using margin (difference between two highest answer probabilities or answer logits) as the uncertainty metric instead of maximum. Left: probability margin AUROC values. These AUROC values and the correlation strength with Q&A accuracy ($R^{2}=0.95,p<10^{-4}$) are similar to MSP. Right: logit margin AUROC values. These AUROC values and the correlation strength with Q&A accuracy ($R^{2}=0.94,p<10^{-4}$) for logit margin are significantly higher than for the maximum logit. Overall, for both probabilities and logits, the same qualitative trends hold, whether the margin or the maximum is used.
Appendix D Post-hoc calibration
As discussed in Section˜2, there exist many methods for rescaling the MSPs of a miscalibrated classifier to improve calibration. One method is Platt scaling (Platt, 2000). Given an MSP $x$, the Platt-scaled MSP is
$\text{Platt}_{A,B}(x)=\frac{1}{1+\exp(Ax+B)}$
where $A$ and $B$ are real-valued parameters learned from data. For training data, we use the same set of $k=20$ data points per dataset as in the Q&A-with-abstention experiments (Section˜6). For each model, we aggregate all the training data across datasets and optimize $A$ and $B$ across the resulting dataset. Figure˜17 and Table˜16 present calibration analysis for the Platt-scaled MSPs. Compared to Figure˜2, calibration error is much less. However, there remains no correlation between Q&A accuracy and calibration error ($R^{2}=0.13,p=0.18$).
Figure 17: Calibration analysis for Platt-scaled MSPs (for chat models and zero-shot prompting). In the left plot, AA ranges from 1.4 to 29.3 with a median of 4.1, while $B$ ranges from -28.3 to -0.5 with a median of -3.1. (The right plot has different values of $A$ and $B$ since as before, we computes calibration error separately for each combination of model, prompt phrasing, and dataset and then average to obtain the calibration error per model.)
Table 16: Calibration error after Platt scaling and Q&A accuracy for each chat model.
LLM Q&A accuracy Calibration error ($\times$100)
Falcon 7B 30.1 6.5
Falcon 40B 40.5 10.5
Llama 2 7B 39.7 13.6
Llama 2 70B 58.7 13.9
Llama 3.0 8B 59.1 12.3
Llama 3.0 70B 78.8 13.2
Llama 3.1 8B 62.0 9.1
Llama 3.1 70B 80.6 11.1
Mistral 7B 57.0 12.3
Mixtral 8x7B 69.9 13.8
SOLAR 10.7B 67.5 15.0
Yi 6B 49.6 12.2
Yi 34B 69.1 11.3
GPT-3.5 Turbo 67.5 13.3
GPT-4o 86.9 11.2
We also explored other parameters for $A$ and $B$. For some choices of $A$ and $B$, we can recover the correlation between calibration error and Q&A accuracy. However, this correlation reverses for other choices of $A$ and $B$. It is perhaps unsurprising that we can obtain arbitrary correlations by manually tweaking these parameters.
Figure 18: Calibration analysis for Platt-scaled MSPs with alternative $A,B$ parameters (for chat models and zero-shot prompting). When we explicitly set $A=2.5$ and $B=0$ (left), we recover a strong negative correlation between calibration error and Q&A accuracy ($R^{2}=0.93,p<10^{-4}$). However, for A=0.5A=-0.5 and $B=0$ (right), this correlation reverses ($R^{2}=0.89,p<10^{-4}$). In both cases, the calibration errors themselves are much larger overall than when $A,B$ are chosen by optimizing over the training data (Figure˜17).
Appendix E Dataset-level results
This section presents per-dataset versions of Figure˜2 (calibration), Table˜2 and Figure˜3 (AUROC), and Table˜4 and Figure˜10 (Q&A with abstention).
E.1 ARC-Challenge
Figure 19: Calibration results for ARC-Challenge. The coefficient of determination between Q&A accuracy and calibration error was $R^{2}=0.21$ ($p=0.08)$.
Table 17: AUROC results for ARC-Challenge. See Table 2 for explanation of the $p$-values.
@@ E.1 ARC-Challenge
Falcon 7B 27.1 52.1 0/2 51.3 0/2
Falcon 40B 47.8 68.7 2/2 67.1 2/2
Llama 2 7B 43.5 63.6 2/2 62.2 2/2
Llama 2 70B 73.7 78.7 2/2 74.3 2/2
Llama 3.0 8B 75.8 81.5 2/2 79.4 2/2
Llama 3.0 70B 92.5 89.3 2/2 77.6 2/2
Llama 3.1 8B 76.6 82.5 2/2 77.6 2/2
Llama 3.1 70B 93.2 90.7 2/2 79.3 2/2
Mistral 7B 70.7 72.1 2/2 68.9 2/2
Mixtral 8x7B 83.4 78.8 2/2 65.9 2/2
SOLAR 10.7B 79.5 76.8 2/2 71.1 2/2
Yi 6B 65.8 75.1 2/2 60.8 2/2
Yi 34B 82.4 84.9 2/2 74.2 2/2
GPT-3.5 Turbo 83.6 84.5 2/2
GPT-4o 96.3 91.9 2/2
Figure 20: AUROC results for ARC-Challenge for MSP (left) and Max Logit (right). Both MSP AUROC and Max Logit AUROC exhibit significant correlations with Q&A accuracy: $R^{2}=0.93$ ($p<10^{-4}$) and $R^{2}=0.68$ ($p<10^{-3}$), respectively.
Table 18: Q&A-with-abstention results for ARC-Challenge. See Table 4 for an explanation of the scoring scheme.
Balanced Conservative
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 45.7-45.7 3.6-3.6 1.6-1.6 118.5-118.5 9.8-9.8 4.3-4.3
Falcon 40B 4.3-4.3 $10.3$ $6.2$ 56.4-56.4 $6.0$ $0.3$
Llama 2 7B 12.9-12.9 $2.9$ $3.7$ 69.4-69.4 0.7-0.7 $1.3$
Llama 2 70B $47.4$ $48.4$ $49.6$ $21.2$ $39.7$ $34.0$
Llama 3.0 8B $51.5$ $53.7$ $52.1$ $27.2$ $43.9$ $39.8$
Llama 3.0 70B $85.0$ $85.0$ $85.0$ $77.5$ $77.5$ $77.5$
Llama 3.1 8B $53.2$ $54.2$ $52.9$ $29.8$ $40.4$ $39.5$
Llama 3.1 70B $86.4$ $86.7$ $86.4$ $79.7$ $81.5$ $79.7$
Mistral 7B $41.6$ $44.4$ $37.5$ $12.3$ $25.9$ $23.7$
Mixtral 8x7B $66.9$ $67.8$ $41.7$ $50.3$ $53.6$ $36.1$
SOLAR 10.7B $59.0$ $59.9$ $59.0$ $38.5$ $43.6$ $43.0$
Yi 6B $31.6$ $34.2$ $31.6$ 2.7-2.7 $24.2$ $9.3$
Yi 34B $64.7$ $65.4$ $59.9$ $47.1$ $49.4$ $50.1$
GPT-3.5 Turbo $67.1$ $67.7$ $50.7$ $58.4$
GPT-4o $92.6$ $92.6$ $88.8$ $88.8$
Figure 21: Q&A-with-abstention scores for ARC-Challenge as a function of the amount of training data. The x-axis is the number of data points included in the training data (referred to as $k$ in Section˜6.1).
E.2 HellaSwag
Figure 22: Calibration results for HellaSwag. The coefficient of determination between Q&A accuracy and calibration error was $R^{2}=0.23$ ($p=0.07)$.
Table 19: AUROC results for HellaSwag. See Table 2 for an explanation of the $p$-values.